OpenStudy (anonymous):

How many molecules of H2O are there in 1.0 g of H2O?

OpenStudy (anonymous):

in 1.0 grams of H2O. okay so.. You are given grams (mass). To go from mass to molecules, you have to convert: Mass -> Moles -> Molecules. To go from mass to moles, Divide by the molecular weight. To go from moles to molecules, multiply by 6.02 x 10^23.

OpenStudy (abb0t):

convert to moles. then use avogradro's number \(6.022 \times 10^{23}\frac{ atoms }{ mol }\)

OpenStudy (anonymous):

I do not know how to do this ..

OpenStudy (anonymous):

Sure you do ;) Let's take it step by step. You are given 1 gram of H2O, and need to divide this by the molecular weight. The molecular weight can be found by adding up all the individual masses of the atoms. How many atoms of Hydrogen are in H2O? How many atoms of Oxygen are in H2O? Multiply the number of atoms of Hydrogen by the atomic mass of hydrogen. Multiply the number of atoms of Oxygen by the atomic mass of oxygen. Add these two values together, what do you get?

OpenStudy (anonymous):

2 atoms in Hydrogen and 1 atom of Oxygen. atomic mass of hydrogen is 1 2*1=1 atomic mass of oxygen is 15.9994 15.994*1=15.9994 Added together is 16.9994

OpenStudy (mertsj):

Do you have it now? Or do you still need help?

OpenStudy (anonymous):

I don't know if 16.9994 is the answer?

OpenStudy (mertsj):

Hang on. I'll figure it out

OpenStudy (anonymous):

Okay thank you

OpenStudy (mertsj):

No. Obviously no. It's way too small.

OpenStudy (anonymous):

I have no clue what I am doing

OpenStudy (mertsj):

Do you know what a mole is?

OpenStudy (anonymous):

Not really

OpenStudy (mertsj):

That might be your trouble. A mole of any compound or element is the molecular weight in grams. So if you had a scale and wanted to weigh out a mole of water, you would weigh out 18 g because 2H=2, O = 16 and the molecular weight is 18

OpenStudy (mertsj):

Do you understand that?

OpenStudy (anonymous):

No, I am so sorry

OpenStudy (mertsj):

Do you know how to find molecular weight?

OpenStudy (anonymous):

I looked it up online

OpenStudy (mertsj):

So I take that to be a yes?

OpenStudy (mertsj):

Do you have a periodic table?

OpenStudy (anonymous):

I looked up the weight online. and I do, yes

OpenStudy (mertsj):

Find the molecular weight of HCl

OpenStudy (anonymous):

36.46094

OpenStudy (mertsj):

So a mole of HCl is 36.46094 grams.

OpenStudy (mertsj):

Now here is the next important point: a mole of "stuff" has 6.023 x 10^23 molecules in it.

OpenStudy (mertsj):

That's Avagadro's number. That is something you cannot reason out you have to learn it.

OpenStudy (mertsj):

So lets say I have some carbon. A mole of carbon is 12 grams. I have 24 grams. How many moles is that?

OpenStudy (anonymous):

12

OpenStudy (anonymous):

2

OpenStudy (mertsj):

Yes. Good. Now remember each mole is 6.023 x 10^23 molecules so how many molecules of carbon do I have?

OpenStudy (anonymous):

I suppose I do not understand how 6.023 X 10^23 works because if there are 2 moles how exactly do I work that in?

OpenStudy (anonymous):

If you have 1 mole of "stuff", you have exactly 6.023 x 10^23 molecules. If you have 2 moles of stuff, you would multiply Avogadro's number by 2.

OpenStudy (mertsj):

Would you agree that 2 moles is twice as much as 1 mole and so I should multiply by 2?

OpenStudy (anonymous):

^

OpenStudy (mertsj):

And \[2(6.023 \times 10^{23)}=12.046 \times 10^{23}\]

OpenStudy (mertsj):

So that is the number of molecules in 2 moles of carbon.

OpenStudy (mertsj):

Ok. dimensionx will help you now. Too many cooks.

OpenStudy (anonymous):

so we aren't really solving the entire equation because when I did that I had a ginormous number

OpenStudy (anonymous):

@Mertsj sorry didn't mean to tread on you xD blahhhh, We are taking it step-by-step ;)

OpenStudy (anonymous):

So we know that we have 2 Hydrogens and 1 Oxygen. 2 x Hydrogen 1 x Oxygen 2 x (1) = 2 1 x (16) = 16 Total Molecular Weight = 18

OpenStudy (anonymous):

Do you understand how I got this value?

OpenStudy (anonymous):

yes.

OpenStudy (mertsj):

So why isn't this problem done yet?

OpenStudy (anonymous):

because I do know how to apply this... I am thinking at 18 is the answer. I am sorry It is just not clicking.

OpenStudy (mertsj):

18 grams is 1 mole of water. How many moles is 9 grams of water?

OpenStudy (anonymous):

.5

OpenStudy (mertsj):

That would be 9/18 right?

OpenStudy (anonymous):

yes

OpenStudy (mertsj):

How many moles is 3 grams of water?

OpenStudy (mertsj):

Just tell me the fraction.

OpenStudy (anonymous):

50 g

OpenStudy (mertsj):

3/18

OpenStudy (mertsj):

I think I see a pattern here: I will draw it because I want you to see it too.

OpenStudy (anonymous):

okay!

OpenStudy (mertsj):